OpenStudy (anonymous):

the graph of 8y+3x=0 is: 1. a line that is neither horizontal nor vertical 2. a vertical line 2. a horizontal line 3. not a line

OpenStudy (raffle_snaffle):

solve for y

OpenStudy (solomonzelman):

it goes through (0,0) right? plug in x=1, and tell me what point you get.

OpenStudy (solomonzelman):

then you should be able to tell right away if this line is horizontal, vertical or neither (based on which points it goes though).

OpenStudy (solomonzelman):

Or, solve for y, as raffle_snaffle advised.

OpenStudy (anonymous):

"Write an equation of the line that passes through the given point and is parallel to the graph of the given equation." #1. (3,2); y=3x-2 #2. (-4,-1); y=2x+14 #3. (-8,6); y=-1/4x+5 #4. (6,2); y=2/3x+19 #5. (10,-5); y=3/2x-7 #6. (-3,4); y=2 "Determine whether the graphs of the given equations are parallel, perpendicular, or neither. Explain." #7. y=4x+5 -4x+y=-13 #8. y=7/9x-7 y=-7/9x+3 #9. y=7/8 x=-4 #10. y=-6x-8 -x+6y=12 #11. 3x+6y=12 y-4=-1/2(x+2) #12. y=4x+12 x+4y=32 "Determine whether each statement is always, sometimes, or never true. Explain." #13. Two lines with different slopes are perpendicular. #14. The slopes of vertical lines and horizontal lines are opposite reciprocals. #15. A vertical line is perpendicular to the x-axis.

OpenStudy (anonymous):

It would be 2. a horizontal line, right?

OpenStudy (anonymous):

yep